# On Safety and Liveness Filtering Using Hamilton-Jacobi Reachability Analysis

### v1

@@ VIII Conclusion And Future Work
In this paper, we explicitly characterize the set of all live and safe control inputs for a dynamical system using Hamilton-Jacobi reachability analysis. Leveraging these maximal sets, we introduce a general framework for designing safety and liveness filters for the system. We propose three distinct safety/liveness filters, balancing performance, control smoothness, and latency, allowing users to select based on system needs.
While these filters integrate safety/liveness and performance, they modify the nominal controller only at the current timestep, unaware of the long-term effects of its action. This might result in myopic controllers. Furthermore, the discrete time nature of actual hardware implementations of these controllers might pose a challenge to the guarantees and analysis presented in this paper. We aim to tackle these challenges in future works.

### v2

@@ VIII Conclusion And Future Work
In this paper, we explicitly characterize the set of all live and safe controls for a dynamical system using Hamilton-Jacobi reachability analysis. Leveraging these maximal sets, we introduce a general framework for designing safety and liveness filters for the system. We propose three distinct safety/liveness filters, balancing performance, control smoothness, and latency, allowing to select based on system needs.
While these filters integrate safety/liveness and performance, they modify the nominal controller only at the current timestep, unaware of the long-term effects of its action. This might result in myopic controllers. Second, the discrete time nature of actual hardware implementations of these controllers might pose a challenge to the guarantees proposed. Third, we currently assume that the target set is known and does not change online. However, that may not be the case for many robotics applications and an online adaptation of the reachability analysis and safety filters might be required. We aim to tackle these challenges in future works. Finally, we currently rely on the differentiability of the value function to construct safety and liveness filters. A theoretical analysis of filter design using subgradients or one-sided gradients of the value function is another important future research direction.